OpenStudy (anonymous):

antiderivative of (x^3)^1/2 + (x^2)^1/3

OpenStudy (anonymous):

\[(x^a)^b = x^{ab}\]\[\int\limits_{}^{}x^n dx = \frac{ x^{n+1} }{ n+1 } + C\]

OpenStudy (anonymous):

is this helpful?